# Transforming faces into video stories - VideoFace2.0 thanks: 4 equal contribution

### v1

@@ Abstract
Face detection and face recognition have been in the focus of vision community since the very beginnings. Inspired by the success of the original Videoface digitizer, a pioneering device that allowed users to capture video signals from any source, we have designed an advanced video analytics tool to efficiently create structured video stories, i.e. identity-based information catalogs. VideoFace2.0 is the name of the developed system for spatial and temporal localization of each unique face in the input video, i.e. face re-identification (ReID), which also allows their cataloging, characterization and creation of structured video outputs for later downstream tasks. Developed near real-time solution is primarily designed to be utilized in application scenarios involving TV production, media analysis, and as an efficient tool for creating large video datasets necessary for training machine learning (ML) models in challenging vision tasks such as lip reading and multimodal speech recognition. Conducted experiments confirm applicability of the proposed face ReID algorithm that is combining the concepts of face detection, face recognition and passive tracking-by-detection in order to achieve robust and efficient face ReID. The system is envisioned as a compact and modular extensions of the existing video production equipment. We hope that the presented work and shared code will stimulate further interest in development of similar, application specific video analysis tools, and lower the entry barrier for production of high-quality multi-modal ML datasets in the future.
@@ I Introduction
Input video from recorded file, network stream or camera feed is uploaded into the system and analyzed frame-by frame in order to detect human faces (spatial and temporal localization), catalog their identity (create a new entry, i.e. video story or re-identify the person if it was already present in the video before), and generate additional personal characteristics with specific time stamps based on further processing of face images and accompanying data (e.g. facial landmarks, facial dynamics, age, gender and audio/speech). Depending on the amount of necessary analyses and available equipment, the system is designed to operate in near real-time mode and allows for standalone embedded hardware implementation.
@@
Another important application field includes production of custom video-based datasets that are required in different ML tasks involving multi-modal inputs like speech, text and image. Challenging vision tasks like lip reading and multi-modal speech recognition require large-scale language and speaker specific video datasets necessary for ML model training. Thus, one of the main goals of the proposed functionality, designed elements and novel algorithmic solutions for face re-identification in VideoFace2.0 was also to facilitate creation of high-quality multi-modal datasets involving human speakers.
@@ III Method and implementation
2:    Initialization: $I(t)$, ${\cal{G}}$, ${\cal{D}}$, $\sigma_{h}$, ${\cal{R}}$, $d(\cdot,)$, $\tau_{d}$, ${\cal{T}}$, $\tau$, $t_{min}$
@@
In order to further improve ReID performance, additional passive tracker ${\cal{T}}$ was also included in Algorithm 1, however in a careful way that does not increase processing latency. It means that it was designed not to be constantly active, but only when needed, in addition to being computationally light (tracking by detection). Namely, since the detector and recognizer are successfully complementing each other most of the time, the only situations in which the tracker is invoked are when additional confirmation of the detection of new identity is necessary. In such cases, standard Intersection-over-Union (IoU) metric and the chosen IoU threshold $\tau$ are utilized to confirm that new identity has entered the video, step 4.1 in Algorithm 1. We note that such setting was possible since in the planned application scenarios, Fig. 1, faces are expected to enter the scene gradually, and it is highly unlikely that a new face will appear immediately over the same position of some face from the previous frame.
In addition to the described algorithmic elements that introduce processing latency due to their inherent computational complexity or hardware limitations, there is also a final, post filtering step where the deliberately introduced delay $t_{min}$ can be traded for additional protection against fake new identities.
@@ IV Results and discussion
Extensive experiments of the proposed open-set face ReID procedure were conducted based on the code implementation provided in Section I. Experiments included visual assessment of ReID results based on specially created visuals and video stories, statistical analysis of processing log files, tests on video sequences with varying number of people, dynamics and shooting conditions, Fig. 2. The results of these experiments and three selected test sequences are available in the code repository. They also include side-by-side visual comparisons of the ablation experiments, Fig. 2(d), in which the impact of each of the described algorithmic components on the ReID performance was analyzed. We note that the given implementation is modular and independent of the specific choice of methods for each of the components in Algorithm 1.
@@
Regarding the implementation we would like to highlight the following capabilities. First, the ability to choose between near real-time generation of specified output video or efficient off-line post production of different types of video output based on detailed log of initial video analysis that is conducted only once. The log file contains information about the temporal and spatial localization of each valid identity in the gallery, their face embeddings and, depending on the settings, other data (gender and age estimates, distinctive face landmarks for each detection, and other statistics, Fig. 2(e)). Control flags and parameters allow fine-tuning of functions and their operation. Predefined video stories, with and without sound, for the user-selected identity include: a) ReID visualizations (overlays of various processing results over original video), Fig. 2(d); b) cropped and optionally scaled face videos, Fig. 2(a), or c) mouth region videos, Fig. 2(b). Note that generated log files contain only anonymized data (information about face appearances), and as such are suitable for sharing among the research community even if the original videos do not allow sharing without specific licenses.
@@
Experimental setup was based on SCRFD face detector [15] and ArcFace embeddings [16], which have been trained on the subset of WebFace dataset [17] and kindly provided by InsightFace project [18], i.e. buffalo_l model pack[19], with overall model size of 325 MB. VideoFace2.0 was tested on Intel i-7 notebook with 16GB of RAM and 4GB NVIDIA RTX3050 GPU. For efficient inference we have utilized ONNX runtime [20] with CUDA support, instead of possibly more efficient TensorRT [21], and achieved a processing speed between 18-25 fps. Relatively small model size should allow for implementation on embedded platforms like Jetson Nano[22], which is planned future work. For image processing and video coding we have relied on OpenCV [23] and FFmpeg [24] libraries. In all experiments, in case of full Algorithm 1, lower-right image in Fig. 2(d), we used: $\sigma_{h}=0.6$, ReID similarity score threshold of $0.4$, i.e. $\tau_{d}=0.6$, $\tau=0.8$, and $t_{min}=60$ frames. Introduced $t_{min}$ delay in new identity approval only affects new identities, and could be replaced by more complex ReID decision rule.

### v2

@@ Abstract
Face detection and face recognition have been in the focus of vision community since the very beginnings. Inspired by the success of the original Videoface digitizer, a pioneering device that allowed users to capture video signals from any source, we have designed an advanced video analytics tool to efficiently create structured video stories, i.e. identity-based information catalogs. VideoFace2.0 is the name of the developed system for spatial and temporal localization of each unique face in the input video, i.e. face re-identification (ReID), which also allows their cataloging, characterization and creation of structured video outputs for later downstream tasks. Developed near real-time solution is primarily designed to be utilized in application scenarios involving TV production, media analysis, and as an efficient tool for creating large video datasets necessary for training machine learning (ML) models in challenging vision tasks such as lip reading and multimodal speech recognition. Conducted experiments confirm applicability of the proposed face ReID algorithm that is combining the concepts of face detection, face recognition and passive tracking-by-detection in order to achieve robust and efficient face ReID. The system is envisioned as a compact and modular extensions of the existing video production equipment. Presented results are based on test implementation that achieves between 18–25 fps on consumer type notebook. Ablation experiments also confirmed that the proposed algorithm brings relative gain in the reduction of number of false identities in the range of 73%–93%.
@@ I Introduction
Input video from recorded file, network stream or camera feed is uploaded into the system and analyzed frame-by frame in order to detect human faces (spatial and temporal localization), catalog their identity (create a new entry, i.e. video story or re-identify the person if it was already present in the video before). It can also generate additional personal characteristics with specific time stamps (e.g. facial landmarks, facial dynamics, age, gender and audio/speech). The system is designed to operate in near real-time mode and allows for standalone embedded hardware implementation.
@@
Another important application field includes production of custom video-based datasets that are required in different ML tasks involving multi-modal inputs like speech, text and image. Challenging vision tasks like lip reading and multi-modal speech recognition require large-scale language and speaker specific video datasets necessary for ML model training. Thus, one of the main goals of VideoFace2.0 and the proposed functionality and designed elements was also to facilitate creation of high-quality multi-modal datasets involving human speakers. The main contribution of the paper is the generic and modular algorithm for face ReID, which is combining the concepts of face detection, recognition and passive tracking-by-detection.
@@ III Method and implementation
2:    Initialization: $I(t)$, video frame at time $t$; open-set face gallery ${\cal{G}}$; face detector ${\cal{D}}$ and confidence threshold $\sigma_{h}$; face embedding ${\cal{R}}$; distance $d(\cdot,)$ and threshold $\tau_{d}$; passive tracker ${\cal{T}}$ and IoU threshold $\tau$; post filtering period $t_{min}$;
@@
In order to further improve ReID performance, additional passive tracker ${\cal{T}}$ was also included in Algorithm 1, however in a careful way that does not increase processing latency. It means that it was designed not to be constantly active, but only when needed, in addition to being computationally light (tracking-by-detection). Namely, since the detector and recognizer are successfully complementing each other most of the time, the only situations in which the tracker is invoked are when additional confirmation of the detection of new identity is necessary. In such cases, standard Intersection-over-Union (IoU) metric and the chosen IoU threshold $\tau$ are utilized to confirm that new identity has entered the video, step 4.1 in Algorithm 1. We note that such setting was possible since in the planned application scenarios, Fig. 1, faces are expected to enter the scene gradually, and it is highly unlikely that a new face will appear immediately over the same position of some face from the previous frame.
In addition to the described algorithmic elements, there is also a final, post filtering step where the deliberately introduced delay $t_{min}$ can be traded for additional protection against fake new identities. However, introduced $t_{min}$ delay in new identity approval only affects new identities when they appear for the first time, and could be replaced by a more complex ReID rule.
@@ IV Results and discussion
Extensive experiments of the proposed open-set face ReID procedure were conducted based on the code implementation provided in Section I. Experiments included visual assessment of ReID results on created visuals and video stories, statistical analysis of processing log files, and tests on three representative video sequences with varying number of people, dynamics and shooting conditions, Fig. 2. These also included side-by-side visual comparisons of the ablation experiments, Fig. 2(d), in which the impact of each of the described algorithmic components was analyzed, Table I. For more information about ablation experiments and test sequences please see the descriptions in the provided code repository. We note that the given implementation is modular and allows for independent choice of methods for each of the components in Algorithm 1.
@@
Regarding the implementation we would like to highlight the following capabilities. First, the ability to choose between near real-time generation of specified output video or efficient off-line post production of different types of video output based on detailed log of initial video analysis that is conducted only once. The log file contains information about the temporal and spatial localization of each valid identity in the gallery, their face embeddings and, depending on the settings, other data (gender and age estimates, distinctive face landmarks for each detection, and other statistics, Fig. 2(e)). Control flags and parameters allow fine-tuning of functions and their operation. Predefined video stories, with and without sound, for the user-selected identity include: a) ReID visualizations (overlays of various processing results over original video), Fig. 2(d); b) cropped and optionally scaled face videos, Fig. 2(a), or c) mouth region videos, Fig. 2(b).
TABLE I: Reduction of false identities brought by Algorithm 1 - ablation study
#identities $\exp 1$ $\exp 2$ $\exp 3$ $\exp 4$ true $\gamma$a  [%] [m:s]b
testvideo1 50 42 30 7 4 83 02:44
testvideo2 421 378 263 23 13 93 07:25
testvideo3 39 37 25 9 6 73 18:45
  • a
    Alg. 1 relative gain: $\gamma=(1-\exp 4/(\sum_{i=1}^{3}(\exp i)/3))\times 100\%$; b duration
Note that generated log files contain only anonymized data (information about face appearances), and as such are suitable for sharing even if the original videos require specific licenses.
@@
Experimental setup was based on SCRFD face detector [15] and ArcFace embeddings [16], which have been trained on the subset of WebFace dataset [17] and kindly provided by InsightFace project [18], i.e. buffalo_l model pack[19], with overall model size of 325 MB. VideoFace2.0 was tested on Intel i-7 notebook with 16GB of RAM and 4GB NVIDIA RTX3050 GPU. For efficient inference we have utilized ONNX runtime [20] with CUDA support, instead of possibly more efficient TensorRT [21], and achieved a processing speed between 18-25 fps. Relatively small model size should allow for implementation on embedded platforms like Jetson Nano[22], which is planned future work. For image processing and video coding we have relied on OpenCV [23] and FFmpeg [24] libraries. In all experiments involving full Algorithm 1, lower-right image in Fig. 2(d), parameters were pre-set to $\sigma_{h}=0.6$, ReID similarity score threshold of $0.4$, i.e. $\tau_{d}=0.6$, $\tau=0.8$, and $t_{min}=60$ frames, based on extensive pre-testing. We note that these are dependant on the specific choice of components in Algorithm 1.